OpenStudy (anonymous):

how to solve f(x)=x^2-3x-2

OpenStudy (anonymous):

will u elaborate

OpenStudy (zzr0ck3r):

there is nothing to solve

OpenStudy (zzr0ck3r):

you may solve x^2-3x-2=0 or you may find the zeros of f(x) but solving f(x) makes no sense.

OpenStudy (anonymous):

given the function rule f(x)= x^2 -3x+2, what is the out put of f(-2)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

i need to know how did u do it

OpenStudy (anonymous):

not the awenser

OpenStudy (zzr0ck3r):

put a -2 where you see an x \(f(x) = x^2-3x-2\\f(-2) = (-2)^2-3(-2)-2 = 4+6-2=8\)

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!

OpenStudy (anonymous):

-2^2 -3(-2)+2=12 @zzr0ck3r

OpenStudy (zzr0ck3r):

its not -2^2 its (-2)^2

OpenStudy (zzr0ck3r):

-2*-2 = 4

OpenStudy (zzr0ck3r):

one must be careful when plugging in cvalues, I always use the () as I did ablove as not to confuse myself...

OpenStudy (zzr0ck3r):

values*

OpenStudy (zzr0ck3r):

@jaga_co understand?

OpenStudy (anonymous):

so who is right?

OpenStudy (anonymous):

before give ans see the problem

OpenStudy (anonymous):

given the function rule f(x)= x^2 -3x+2, what is the out put of f(-2) @zzr0ck3r

OpenStudy (zzr0ck3r):

I am right. I promise

OpenStudy (anonymous):

i know for a fackt that the awense is 12 but i dont know how to work it out

OpenStudy (anonymous):

fact*

OpenStudy (zzr0ck3r):

that topic says -2

OpenStudy (zzr0ck3r):

f(x)=x^2-3x-2

OpenStudy (zzr0ck3r):

I am confused

OpenStudy (zzr0ck3r):

OHHHHHHHH lol you changed the equation

OpenStudy (zzr0ck3r):

sorry

OpenStudy (anonymous):

thanks @zzr0ck3r for your confusion

OpenStudy (zzr0ck3r):

@jaga_co is correct

OpenStudy (anonymous):

lol its kk :)

OpenStudy (zzr0ck3r):

notice the thing in the topic is not the same thing as in the follow up

OpenStudy (anonymous):

i wish i could give u boath a medal but i have to give it to jaga_co :)

OpenStudy (zzr0ck3r):

lol I dont need medals:)

OpenStudy (zzr0ck3r):

I medal-ed him also:)

OpenStudy (anonymous):

kk :) thanks aney way